# Bayesian Ensembling: Insights from Online Optimization and Empirical Bayes

### v1

@@ 3.2 Online Bayesian Stacking is Portfolio Selection
A point of nuance in applying OPS algorithms is the existence of a market variability parameter $\alpha$, requiring pre-determined maximum and minimum predictive densities. The maximum is clear for proper Bayesian models, and a minimum may be assumed (e.g., via compact data spaces, or bounded model error). However, $\alpha$ may still be very small, which can be problematic for EG/ONS regret analysis. Using subgradient norms in regret analysis can bound tighter bounds, as non-adversarial predictions from failing models (i.e., $w_{t,k}\approx 0$) are unlikely to suddenly dominate. If such outliers are a concern, Soft-Bayes also provides $\alpha$-independent regret bounds.
We provide pseudocode for OBS and O-BMA in Algorithm 1. The pseudocode underscores the similarity between OBS and O-BMA, replacing a single update equation. Other algorithms for OPS may be used by changing Line 6.
3.3 Online Bayesian Model Averaging is the Hedge Algorithm with Learning Rate $1$
Now that we have established an equivalence between OBS and the portfolio selection problem, one may wonder if a similar connection holds for O-BMA. Indeed, it is classically known that O-BMA updates are a specific choice of the Hedge algorith (Freund & Schapire, 1997), which performs expert selection. While this connection is established (Vovk, 2001), we rederive it as a useful narrative tool to motivate O-BMA as a model selection algorithm, which we show clearly using Lemma 3.1.
@@ 3.3 Online Bayesian Model Averaging is the Hedge Algorithm with Learning Rate $1$
whose gradient is $\nabla_{w_{t,k}}\ell_{t}(\boldsymbol{\mathbf{w_{t}}})=-\log p_{k}(y_{t}|\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})$. This suggests that O-BMA may act similar to an OCO algorithm that minimizes regret with respect to a different loss function,
@@
This reveals a key insight: OBS aims for the best post-hoc expert mixture, while O-BMA targets the best single expert. The following lemma (which is a known result, but presented in an unconventional form), proved in Section B.2, shows how O-BMA performs compared to the best expert.
Lemma 3.1.
@@
Typically, the regret in Lemma 3.1 is simplified by taking uniform priors $\Pr(\mathcal{M}_{k})=1/K$ and noting that $\log\Pr(\mathcal{M}_{k^{*}}|\mathcal{D}_{T})<0$, so that
$R_{T}<\log K.$ (15)
This shows BMA has very strong regret guarantees with respect to the best expert, never doing much worse than the best expert.
The typically-presented regret bound Equation 15 does not preclude the situation where BMA performs better than the best expert. However, using the equality in Lemma 3.1, we can also argue that BMA also does not typically do much better than the best expert. In particular, if $\Pr(\mathcal{M}_{k^{*}}\,|\,\mathcal{D})$ is bounded below, so is the regret. As an important special case, if $\Pr(\mathcal{M}_{k^{*}}\,|\,\mathcal{D})>\Pr(\mathcal{M}_{k^{*}})$ — as one might typically expect — the regret is strictly positive, and the predictive log-likelihood of the BMA mixture performs almost exactly as the best model does, on average. Therefore, we may study O-BMA as an online optimizer that minimizes the regret to Equation 13, i.e., an expert selection problem. We use this insight to derive a novel empirical Bayes perspective in the next section.
@@ 3.4 Regret Analysis
We further discuss regret bounds in Appendix B. In Section B.3, we give an example that shows, with mild additional assumptions, we can even recover regret bounds from O-BMA applications, albeit with potentially worse constants — this is despite the fact that O-BMA cannot provide comparable bounds when regret is measured against the mixture loss.
@@ 4 An Empirical Bayes Perspective
An empirical Bayes (EB) perspective (Robbins, 1964) elucidates BMA and Bayesian stacking properties, offering direct arguments for their limitations. Recall that in EB, hyperparameters $\boldsymbol{\mathbf{\psi}}$ are found by maximizing the marginal likelihood, rather than being marginalized out. We show O-BMA and OBS have subtly distinct EB justifications. While we focus notationally on the online setting, similar results hold for batch processing.
@@ 4.1 BMA is Empirical Bayes over an Indicator Variable
It is well-known in the literature that BMA collapses to the model with the highest marginal likelihood when the amount of data increases (Yao et al., 2018), but explanations of this fact are not often presented. Using our insights that O-BMA is the Hedge algorithm, i.e., that it minimizes regret with the loss function (13), this fact is straightforward to see: $w_{k}$ should be $1$ for the model with the highest marginal likelihood and 00 for the others. In this sense, we may interpret BMA as performing empirical Bayes over an indicator variable $\boldsymbol{\mathbf{z}}$ in the corresponding mixture distribution. To see this, we note that (13) can be written as
@@
and that $\sum_{k=1}^{K}w_{k}\log p_{k}(\mathcal{D}_{T})\leq\max_{k}\log p_{k}(\mathcal{D}_{T}).$ Hence, the optimal weight vector corresponds to $w_{k}=0$ for all $k$ except for the model $k^{*}=\operatorname*{arg\,max}_{k}p_{k}(\mathcal{D}_{T})$ with highest marginal likelihood, for which $w_{k^{*}}=1$. Note also that by Jensen’s inequality, (17) is a lower bound to $\log(\sum_{k}w_{k}p_{k}(\mathcal{D}_{T}))$. If we set $w_{k}=\text{Pr}(\mathcal{M}_{k})$, this is the log of the evidence of the BMA model, so we can interpret BMA as performing empirical Bayes on the discrete prior probabilities of the models.
@@ 4.2 OBS is Empirical Bayes Estimation over Mixture Weights
On the other hand, OBS can be seen as performing empirical Bayes of the mixture weights themselves. To see this, we interpret $\boldsymbol{\mathbf{\psi}}$ as $\boldsymbol{\mathbf{w}}$, and the posterior predictive density
@@
Hence, in stacking we obtain the weights by maximizing the log-evidence of the ‘stacking’ model, $\log p_{\text{stack}}(\mathcal{D}_{t}|\boldsymbol{\mathbf{w}})$, where the weights correspond to hyperparameters. Note the difference between this marginal likelihood and the one in BMA, where the marginal likelihood is itself a mixture and each component is independent of $\boldsymbol{\mathbf{w}}$.
@@ 5 Experiments
Thus far, we have primarily focused on the theoretical properties of OBS. In this section, we provide empirical evidence that OBS can be beneficial. We consider four main scenarios: an illustrative toy example (Section 5.1), variational Bayesian neural networks (Section 5.2), SMC-based stochastic volatility models (Section 5.4), and dynamic Gaussian processes in non-stationary environments (Section 5.4). We provide details on our experimental setup, baselines, and code in Appendix C. Throughout, we use a set of default OCO hyperparameters, with some ablation in Appendix H.
(a) Toy Example: Open (Section 5.1)
(b) Toy Example: Closed (Section 5.1)
Figure 1: The average predictive log-likelihood (higher is better) in the toy example. Lines denote the median and shaded area represent the 10th to 90th percentiles over 10 trials. The first 100 samples are suppressed for readability.
@@ 5.1 Subset Linear Regression
We revisit a classical problem from Breiman (1996), also used by Yao et al. (2018), generating i.i.d. data from a 15-dimensional Gaussian linear model with weak predictors as in Yao et al. (2018, Section 4.2). We consider two scenarios with an ensemble of 15 regression models. In the “Open” setting, the ensemble consists of 15 univariate models ($y\approx\theta_{k}x_{k}$), none matching the true process, where performance depends on combining models. In the “Closed” setting, model $k$ uses features $x_{1},\dots,x_{k}$, making model 15 the true model. In both scenarios, hyperparameters were pre-trained using empirical Bayes on 1000 points, followed by online deployment of OBS or O-BMA for 5000 points. Further experimental details are in Appendix D. As shown in Figures 1(a) and 1(b), OBS variants (EG, ONS) dramatically outperform O-BMA in the “open” setting. In the “closed” setting, OBS performs nearly as well as the theoretically optimal O-BMA, with ONS slightly outperforming EG.
@@ 5.2 Online Variational Inference
We now move to a more practical application in Bayesian machine learning, applying OBS to online Bayesian neural networks. We use the recently proposed Bayesian online natural gradient (BONG), which optimizes the expected log-likelihood with online mirror descent (Jones et al., 2024). Many different variants of BONG and related approaches are tested, unifying them under a common framework; we make online predictions on MNIST (LeCun et al., 2010) using the best variant tested, “bong-dlr10-ef_lin”. To create ensembles, we train five Bayesian feedforward neural networks with different initializations. We provide more details on the experimental setup in Appendix E.
We visualize the resulting predictive log-likelihoods in Figure 2(a). Though less dramatic than in the toy example, we observe clear improvements in predictive performance using OBS instead of O-BMA and DMA. Furthermore, the specific OCO algorithm used does not seem to matter much, with Soft-Bayes, EG, and ONS all resulting in solutions similar to the BCRP. In Appendix E, we additionally visualize the evolution and final weights, which show that BMA collapses to a single model, whereas OBS converges to the optimal, more balanced weights.
@@ 5.3 Online Forecasting
(a) Bayesian Neural Network (Section 5.2)
(b) Forecasting with GARCH (Section 5.4)
Figure 2: The average predictive log-likelihood (higher is better) Bayesian neural network and forecasting experiments. Lines denote the median and shaded area represent the 10th to 90th percentiles over 10 trials. The first 100 samples are suppressed for readability.
We compare BMA and OBS variants for forecasting S&P data, used in an offline analysis by Geweke & Amisano (2011, 2012), ensembling diverse GARCH models. Model parameters are estimated online using sequential Monte Carlo (Doucet et al., 2001), where the typical “bank of filters” approach is equivalent to BMA. Eight models (2 per class, with priors sampled from uniform hyperpriors) were generated – see Appendix F for further details.
We visualize the resulting predictive log-likelihoods in Figure 2(b). Again, we observe clear improvements in predictive performance of OBS over O-BMA, with similar results obtained by Soft-Bayes, EG, and ONS. As this is real-world financial data, we expect some amount of non-stationarity: indeed, we find that DMA improves upon BMA, but remains inferior to the OBS methods, and that D-ONS performs the best. We show that BMA collapses once more in this example in Appendix F, which can be seen in the evolution and final weights.

### v2

@@ 3.2 Online Bayesian Stacking is Portfolio Selection
A point of nuance in applying OPS algorithms is the existence of a market variability parameter $\alpha$, requiring pre-determined maximum and minimum predictive densities. The maximum is clear for proper Bayesian models, and a minimum may be assumed (e.g., via compact data spaces or bounded model error). However, $\alpha$ may still be very small, which can be problematic for EG/ONS regret analysis. Bounding regret using subgradient norms can produce tighter guarantees, since non-adversarial predictions from failing models (i.e., $w_{t,k}\approx 0$) are unlikely to suddenly dominate. If such outliers are a concern, Soft-Bayes also provides $\alpha$-independent regret bounds.
Algorithm 1 Online Ensemble Update: OBS with EG or OBS with Soft-Bayes or O-BMA.
2:for $t=1$ to $T$ do
2:   // Prediction Step
4:  Form ensemble prediction via Equation˜1
5:  Output prediction and receive true label $y_{t}$.
5:   // Update Step
6:  Update model weights using OBS w/ EG (10) or OBS w/ Soft-Bayes (11) or O-BMA (3)
8:end for
We provide pseudocode for OBS and O-BMA in Algorithm˜1. The pseudocode underscores the similarity between OBS and O-BMA, replacing a single update equation. Other algorithms for OPS may be used by changing Line 6.
3.3 Online Bayesian Model Averaging is the Hedge Algorithm with Learning Rate 11
Now that we have established an equivalence between OBS and the portfolio selection problem, one may wonder if a similar connection holds for O-BMA. Indeed, it is classically known that O-BMA updates are a specific choice of the Hedge algorithm (Freund & Schapire, 1997), which performs expert selection. While this connection is established (Vovk, 2001), we rederive it as a useful narrative tool to motivate O-BMA as a model selection algorithm, which we show clearly using Proposition˜3.1.
@@ 3.3 Online Bayesian Model Averaging is the Hedge Algorithm with Learning Rate 11
whose gradient is $\nabla_{w_{t,k}}\ell_{t}(\boldsymbol{\mathbf{w_{t}}})=-\log p_{k}(y_{t}|\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})$. This suggests that O-BMA may act similarly to an OCO algorithm that minimizes regret with respect to a different loss function,
@@
This reveals a key insight: OBS aims for the best post-hoc expert mixture, while O-BMA targets the best single expert. The corresponding bound achieves constant regret; in fact, in the typical proof of this regret upper bound, we prove an equality regarding the regret. The following result is therefore known, but is presented in a somewhat unorthodox way to allow us to show lower bounds on the regret.
Proposition 3.1.
@@
Thus, if $\Pr(\mathcal{M}_{k^{*}}\,|\,\mathcal{D})$ is bounded below, so is the regret. In particular, Proposition˜3.1 applied the following theorem directly, which asserts O-BMA acts as an “optimizer” to Equation˜13.
Corollary 3.2.
Let the regret of the BMA mixture with respect to the best individual model be defined as in Proposition˜3.1. If the posterior probability of the optimal model $\mathcal{M}_{k^{*}}$ exceeds its prior probability, i.e., $\Pr(\mathcal{M}_{k^{*}}\,|\,\mathcal{D})\geq\Pr(\mathcal{M}_{k^{*}})$, then $R_{T}$ is bounded both above and below,
$0\leq R_{T}\leq-\log\Pr(\mathcal{M}_{k^{*}}).$ (15)
Thus, under typical scenarios (e.g., uniform priors on $\mathcal{M}_{k}$), Equation˜15 becomes fairly tight, and the solutions become strong online optimizers of Equation˜15.
@@ 3.4 Regret Analysis
We further discuss regret bounds in Appendix˜A. In Section˜A.3, we provide an example that shows that, with mild additional assumptions, we can even recover regret bounds from O-BMA applications, albeit with potentially worse constants - this is despite the fact that O-BMA cannot provide comparable bounds when regret is measured against the mixture loss.
3.5 A Hybrid Approach
In the case where the M-closed scenario is plausible, but this fact is unknown, OBS still achieves vanishing regret with respect to the singular best expert: this follows because the problem of determining an optimal mixture of densities is harder than the corresponding expert selection problem, which is included as a special case.
Nevertheless, as determined in Proposition˜3.1, the corresponding regret guarantees for O-BMA in the M-closed setting are extremely strong, achieving constant regret as an upper bound. When a model extremely close to the data generating process is available, OBS is thus potentially suboptimal.
One potential approach to ameliorate this is to consider a hybrid method, where O-BMA and OBS mixtures are maintained and further averaged via a secondary layer of O-BMA. That is, we consider two sets of weights, $\boldsymbol{\mathbf{w}}_{t}^{\text{O-BMA}}$ and $\boldsymbol{\mathbf{w}}_{t}^{\text{OBS}}$, along with a secondary set of weights $\boldsymbol{\mathbf{v}}_{t}\in\mathbb{S}^{2}$. The corresponding predictive distribution is thus
$p^{\text{hybrid}}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{1:t-1})=v_{1}\underbrace{\left(\sum_{k}w_{t,k}^{\text{O-BMA}}p_{k}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})\right)}_{p^{\text{O-BMA}}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})}+v_{2}\underbrace{\left(\sum_{k}w_{t,k}^{\text{OBS}}p_{k}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})\right)}_{p^{\text{OBS}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})}}.$ (16)
The predictive performance of this mixture achieves constant regret with respect to the best model, once again, while maintaining the expressiveness of the corresponding OBS solution. In particular, we may obtain the following bound:
Theorem 3.3.
Let the regret of the hybrid mixture in Equation˜16 for the best individual model be defined as in Proposition˜3.1, i.e., $R_{T}=\sum_{t}\log p_{k^{*}}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})-\sum_{t}\log p^{\text{hybrid}}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})$, where $k^{*}$ is the model with the largest marginal likelihood. Further assume uniform prior weights over $\boldsymbol{\mathbf{v}}$ and $\boldsymbol{\mathbf{w}}^{\text{O-BMA}}$. Then $R_{T}$ may be bounded as
$R_{T}\leq\log K+\log 2.$ (17)
Proof.
From the “outermost” perspective of ensembling (i.e., with $\boldsymbol{\mathbf{v}}$), we have by Proposition˜3.1 a regret bound with respect to any “outer” expert (i.e., the O-BMA or OBS mixtures):
$\sum_{t}\log p^{\text{O-BMA}}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})-\sum_{t}\log p^{\text{hybrid}}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})\leq\log 2.$ (18)
Further, from Proposition˜3.1, O-BMA achieves a bounded regret
$\sum_{t}\log p_{k^{*}}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})-\sum_{t}\log p^{\text{O-BMA}}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})\leq\log K.$ (19)
Combining Equation˜18 and Equation˜19 completes the theorem. ∎
We note by a symmetric argument that if the OBS mixture has a clearly stronger predictive likelihood, the “outer” level of O-BMA will select the OBS mixture as the best expert, similarly incurring a simple additive regret factor.
@@ 4 An Empirical Bayes Perspective
An empirical Bayes (EB) perspective (Robbins, 1964) elucidates BMA and Bayesian stacking properties, offering direct arguments for their limitations. Recall that in EB, hyperparameters $\boldsymbol{\mathbf{\psi}}$ are found by maximizing the marginal likelihood, rather than being marginalized out. We show that O-BMA and OBS have subtly distinct EB justifications. While we focus notationally on the online setting, similar results hold for batch processing.
@@ 4.1 BMA is Empirical Bayes over an Indicator Variable
It is well-known in the literature that BMA collapses to the model with the highest marginal likelihood when the amount of data increases (Yao et al., 2018), but explanations of this fact are not often presented. Using our insights that O-BMA is the Hedge algorithm, i.e., that it minimizes regret with the loss function (13), this fact is straightforward to see: $w_{k}$ should be $1$ for the model with the highest marginal likelihood and 0 for the others. In this sense, we may interpret BMA as performing empirical Bayes over an indicator variable $\boldsymbol{\mathbf{z}}$ in the corresponding mixture distribution. To see this, we note that (13) can be written as
@@
and that $\sum_{k=1}^{K}w_{k}\log p_{k}(\mathcal{D}_{T})\leq\max_{k}\log p_{k}(\mathcal{D}_{T}).$ Hence, the optimal weight vector corresponds to $w_{k}=0$ for all $k$ except for the model $k^{*}=\operatorname*{arg\,max}_{k}p_{k}(\mathcal{D}_{T})$ with the highest marginal likelihood, for which $w_{k^{*}}=1$. Note also that by Jensen’s inequality, (21) is a lower bound to $\log(\sum_{k}w_{k}p_{k}(\mathcal{D}_{T}))$. If we set $w_{k}=\text{Pr}(\mathcal{M}_{k})$, this is the log of the evidence of the BMA model, so we can interpret BMA as performing empirical Bayes on the discrete prior probabilities of the models.
@@ 4.2 OBS is Empirical Bayes Estimation over Mixture Weights
On the other hand, OBS can be seen as performing empirical Bayes on the mixture weights themselves. To see this, we interpret $\boldsymbol{\mathbf{\psi}}$ as $\boldsymbol{\mathbf{w}}$, and the posterior predictive density
@@
Hence, in stacking, we obtain the weights by maximizing the log-evidence of the ‘stacking’ model, $\log p_{\text{stack}}(\mathcal{D}_{t}|\boldsymbol{\mathbf{w}})$, where the weights correspond to hyperparameters. Note the difference between this marginal likelihood and the one in BMA, where the marginal likelihood is itself a mixture and each component is independent of $\boldsymbol{\mathbf{w}}$.
@@ 5 Experiments
Thus far, we have primarily focused on the theoretical properties of OBS. In this section, we provide empirical evidence that OBS can be beneficial. We consider four main scenarios: an illustrative toy example (Section˜5.1), variational Bayesian neural networks (Section˜5.2), SMC-based stochastic volatility models (Section˜5.3), and dynamic Gaussian processes in non-stationary environments (Section˜5.4). We provide details on our experimental setup, baselines, and code in Appendix˜B.
(a) Toy Example: Open (Section˜5.1)
(b) Toy Example: Closed (Section˜5.1)
Figure 1: The average predictive log-likelihood (higher is better) in the toy example. ‘EG” is exponentiated gradients, “ONS” is the Online Newton Step, “BCRP” is the optimal constant rebalanced portfolio (offline baseline), and “O-BMA” is O-BMA. Lines denote the median and shaded area represent the 10th to 90th percentiles over 10 trials. The first 100 samples are suppressed for readability.
@@ 5.1 Subset Linear Regression
We revisit a classical problem from Breiman (1996), also used by Yao et al. (2018), generating i.i.d. data from a 15-dimensional Gaussian linear model with weak predictors, as in Yao et al. (2018, Section 4.2). We consider two scenarios with an ensemble of 15 regression models. In the “Open” setting, the ensemble consists of 15 univariate models ($y\approx\theta_{k}x_{k}$), none of which match the true process, where performance depends on combining models. In the “Closed” setting, model $k$ uses features $x_{1},\dots,x_{k}$, making model 15 the true model. In both scenarios, hyperparameters were pre-trained using empirical Bayes on 1000 points, followed by online deployment of OBS or O-BMA for 5000 points. Further experimental details are in Appendix˜C. As shown in Figures˜1(a) and 1(b), OBS variants (EG, ONS) dramatically outperform O-BMA in the “open” setting. In the “closed” setting, OBS performs nearly as well as the theoretically optimal O-BMA, with ONS slightly outperforming EG.
As detailed in the theoretical section, one reason we may expect O-BMA to perform worse than OBS is the weight collapse of BMA. We can empirically validate this property in our toy experiment, with the expectation that O-BMA collapses to a single weight, and OBS retains a proper “mix” of models.
In Figure˜2(a), we show the final weights in the “open” subset linear regression experiment for O-BMA, OBS, and the BCRP (i.e., the optimal pooling of Geweke & Amisano (2011), or Bayesian stacking from the prequential principle (Yao et al., 2018)). In Figure˜3(a), we show the evolution of the BMA and OBS weights as new data arrive. Finally, in Figure˜2(b) and Figure˜3(b), we show the analogous final weights and the evolution of weights for the “closed” subset linear regression experiment.
(a) Toy Example: Open
(b) Toy Example: Closed
Figure 2: The final weights in the “open” and “closed” subset linear regression experiment.
(a) Toy Example: Open
(b) Toy Example: Closed
Figure 3: The evolution of the weight vector $\boldsymbol{\mathbf{w}}_{t}$ as a function of $t$ in the “open” and “closed” subset linear regression experiment. Results are shown for a single trial due to the noisy nature of the plots. Dots on the right side of a plot denote the final weights of the BCRP.
@@ 5.2 Online Variational Inference
We now move to a more practical application in Bayesian machine learning, applying OBS to online Bayesian neural networks. We use the recently proposed Bayesian online natural gradient (BONG), which optimizes the expected log-likelihood with online mirror descent (Jones et al., 2024). Many different variants of BONG and related approaches are tested, unifying them under a common framework; we make online predictions on MNIST (LeCun et al., 2010) using the best variant tested, “bong-dlr10-ef_lin”. To create ensembles, we train five Bayesian feedforward neural networks with different initializations. We provide more details on the experimental setup in Appendix˜D.
We visualize the resulting predictive log-likelihoods in Figure˜4(a). Although less dramatic than in the toy example, we observe clear improvements in predictive performance using OBS instead of O-BMA and DMA. Furthermore, the specific OCO algorithm used does not seem to matter much, with Soft-Bayes, EG, and ONS all resulting in solutions similar to the BCRP. In Appendix˜D, we additionally visualize the evolution and final weights, which show that BMA collapses to a single model, whereas OBS converges to the optimal, more balanced weights.
@@ 5.3 Online Forecasting
We compare BMA and OBS variants for forecasting S&P data, which were used in an offline analysis by Geweke & Amisano (2011; 2012), ensembling diverse GARCH models. Model parameters are estimated online using sequential Monte Carlo (Doucet et al., 2001), where the typical “bank of filters” approach is equivalent to BMA. Eight models (2 per class, with priors sampled from uniform hyperpriors) were generated – see Appendix˜E for further details.
Method Median Reward
EG 3.46
BMA 3.44
ONS 3.46
Soft-Bayes 3.31
Table 1: Median reward (predictive log-likelihood; higher is better) at the final time step in the GARCH experiment with $100$ models.
We visualize the resulting predictive log-likelihoods in Figure˜4(b). Again, we observe clear improvements in the predictive performance of OBS over O-BMA, with similar results obtained by Soft-Bayes, EG, and ONS. As this is real-world financial data, we expect some amount of non-stationarity: indeed, we find that DMA improves upon BMA, but remains inferior to the OBS methods, and that D-ONS performs the best. We show that BMA collapses once more in this example in Appendix˜E, which can be seen in the evolution and final weights.
We additionally performed our online forecasting experiments with $25$ models from each class, for a total of $100$ models. Repeating across 3 random seeds, we obtain the results in Table˜1. We used an EG learning rate 10 times larger than that used in previous experiments, guided by the logarithmic increase in the optimal theoretical learning rate as the number of models increases. The results show that ONS and EG still outperform BMA, while Soft-Bayes performs worse. We posit that the issues with Soft-Bayes are due to the relatively short time horizon (on the order of $1000$ time steps) relative to the number of models – we are fundamentally trading off some fast adaptation for robustness when obtaining the gradient bound-free regret bounds of Soft-Bayes.
(a) Bayesian Neural Network (Section˜5.2)
(b) Forecasting with GARCH (Section˜5.3)
Figure 4: The average predictive log-likelihood (higher is better) in the MNIST and forecasting experiments, respectively. The method descriptions follow those in Figure˜1, with the addition of Soft-Bayes. Lines denote the median and shaded area represents the 10th to 90th percentiles over 10 trials. The first 100 samples are suppressed for readability.